Pioneer Films has just finished a production of Sharknado 5, the latest action film directed by John Favreau and starring Ian Ziering, Tara Reid and George Clooney. The total production cost was $5million. All the production personnel and actors on Sharknado 5 received a fixed salary (included in the $5 million) and will have no "residual" (equity interest) in the revenues or operating income from the movie. Daniels Communications will handle the marketing of Sharknado 5. Daniels agrees to invest a minimum of $3 million of its own money in marketing the movie and will be paid 20% of the revenues Pioneer itself receives from the box-office receipts. Pioneer receives 62.5% of the total box-office receipts (out of which comes the 20% payment to Daniels).

1. What is the breakeven point to Pioneer for Sharknado 5 expressed in terms of (a) revenues received by Pioneer and (b) total box-office receipts?

2. Assume in its first year of release, the box-office receipts for Sharknado 5 total $300 million. What is the operating income to Pioneer from the movie in its first year?

A year goes by and Pioneer is negotiating for Sharknado 6, the thrilling sequel to the mega-blockbuster Sharknado 5. The negotiation is proving more difficult than the original movie. The budgeted production cost (excluding payments to the director Favreau and the stars Reid and Ziering) for Sharknado 6 is $21 million. The agent negotiating for Favreau, Reid and Ziering proposes either of two contracts:

Contract A: Fixed-salary component of $15 million for Favreau, Ziering and Reid (combined) with no residual interest in the revenues from Sharknado 6.

Contract B: Fixed-salary component of $3million for Favreau, Ziering and Reid (combined) plus a residual of 15% of the revenues Pioneer received from Sharknado 6.

Daniels will market Sharknado 6. It agrees to invest a minimum of $10 million of its own money. Because of its major role in the success of Sharknado 5, Daniels will now be paid 25% of the revenues Pioneer received from the total box-office receipts. Pioneer receives 62.5% of the total box-office receipts (out of which comes the 25% payment to Daniels).

 3. What is the breakeven point for Pioneer expressed in terms of (a) revenues received by Pioneer and (b) total box-office receipts for Sharknado 6 for CONTRACTS A AND B? Explain the difference between the breakeven points for contracts A and B.

4. Assume Sharknado 6 achieves the same $300 million in box-office revenues as Sharknado 5. What is the operating income to Pioneer from Sharknado 6 if it accepts Contract B? Discuss the difference in operating income between the two films.

Answers/Work I have:

1)

For Breakeven, the total inflow shall be $5,000,000 (Cash contributed by Pioneer)

revenue received is 1-0.20 or 80% (since we are only looking at pioneer (1))

Total revenue = 5 million/0.80=$6.25 million or $6,250,000

Total box office receipts are :B*0.625=$6,250,000

B=10 million or $10,000,000

2)

For 300 million = 300*0.625*0.80=$150 million

I am stuck on getting started on 3 and 4, maybe it is the wording?
Answer #1

3) Breakeven Point

For Option A:

For Break even, the total inflow shall be $($21,000,000+ $15,000,000) (Cash contributed by Pioneer)

revenue received is 1-0.25 or 75% (since we are only looking at pioneer (1))

Total revenue = 36 million/0.75=$48 million or $ 48,000,000

Total box office receipts are :B*0.625=$48,000,000

B=76.80 million or $ 76,800,000

For Option B:

For Break even, the total inflow shall be $($21,000,000+ $3,000,000) (Cash contributed by Pioneer)

revenue received is 1-0.25-0.15 or 60% (since we are only looking at pioneer (1))

Total revenue = 24 million/0.60=$40 million or $ 40,000,000

Total box office receipts are :B*0.625=$40,000,000

B=64.00 million or $ 64,000,000

Breakeven Point is better For Option B because Fixed Cost is less in option B in compare to Option A

4. Operating Income = For 300 million = 300*0.625*0.60=$ 112.50 million

Operating Income is less in Sharknado 6 in comparision to Sharknado 5 due to variable cost is increased and total revenue remains the same.